OpenStudy (anonymous):

Solve for x. log(3)*(x-5)+log(3)*(x+3)=2

OpenStudy (anonymous):

same story as the previous one , but I'm not sure about the ln part ^_^

OpenStudy (anonymous):

Just to test my understanding. (x-5)(x-3)=2 x^2-5x+3x-15=2 rearrange it x(x-2)=17 so... x= -2, 17